OpenStudy (anonymous):

Consider a solenoid (n turns/unitlength, radius R) carrying a current that follows \[I(t)=kt\] k constant. Calculate the Poynting vector and use it to show that the flow of energy (per unit length) into the volume occupied by the solenoid is given by, \[\frac{\mathrm{d}W}{\mathrm{d}t}=\frac{\mathrm{d}{\mathrm{d}t}(\frac{1}{2}LI^2)\] where L is the self-inductance per unit length of the solenoid.

OpenStudy (anonymous):

Sorry that should be, \[\frac{\mathrm{d}W}{\mathrm{d}t}=\frac{\mathrm{d}}{\mathrm{d}t}(\frac{1}{2}LI^2)\]

OpenStudy (anonymous):

Hm when I go through I get \[P=(\mu_0 n \dot{I} r/2)(\mu_0 n I)=\mu_0^2...\] This then follows through and we have extra mu's in the final statment. the mu squared seems to come from us deriving E from B which carries extra mu. NVM I realized I wasn't using H in calculating P.